OpenStudy (anonymous):

A student increases the temperature of a 556 cm3 balloon from 278 K to 231 K. Assuming constant pressure, what should the new volume of the view be?

OpenStudy (abmon98):

hink of it this way. As the temperature of the gas increases, the gas molecules will begin to move around more quickly and hit the walls of their container with more force—thus the volume will increase. Use Charles Law. \[V1/V2=T1/T2 \]

OpenStudy (anonymous):

OK I know how to do that. I would do it like this right? P1*V1 P2*V2 (278K)(556Ccm3) =(231K) 154568Kcm3 =(231k) and then that's where I et stuck at.........

OpenStudy (abmon98):

556*231/278=462 cm^3

OpenStudy (abmon98):

Why P1*V1=P2*V2 Whats involved here is temperature and volume not pressure, pressure is constant @Zakiyyah

OpenStudy (abmon98):

P1*V1=P2*V2 is used if temperature is constant

OpenStudy (anonymous):

ok thank you. I just had a little misunderstanding. @Abmon98

OpenStudy (abmon98):

your welcome, if you are still unclear about gas laws check this out http://www.chemguide.co.uk/physical/kt/otherlaws.html